I need to fine the first five terms of the sequence with the rule an=3+(n-1)1/2
Answer by stanbon(75887)

I need to fine the first five terms of
the sequence with the rule
an=3+(n-1)1/2
----------------------------------
a(1) = 3
a(2) = 3 + 1/2 = 3 1/2
a(3) = 3 + 1 = 4
a(4) = 3 + 3/2 = 4 1/2
a(5) = 3 + 2 = 5
==========================================
Cheers,
Stan H.
=============